OpenStudy (anonymous):

is this correct? if not please expalin why Problem => (1/7x^2 – 2/5)(1/2x + ½) My answer is: (1/7X^2 – 2/5)(1/2x + ½) = (1/7 * x^2 – 2/5) * (1/2*x + ½) = 0.071 * x^3 + 0.071 * x^2 – 0.2 * x -0.2 = 0.071 * (x + 1.673) * (x + 1) * x – 1.673

OpenStudy (pfenn1):

Is this the original expression?\[\frac{1}{7x^2} – \frac{2}{5}\left( \frac{1}{2x} +\frac 12 \right)\]

OpenStudy (anonymous):

I am lost

OpenStudy (pfenn1):

I think I made a mistake above and the expression should be this.\[\left( \frac{1}{7x^2} – \frac{2}{5} \right)\left( \frac{1}{2x} +\frac 12 \right)\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

Well, I am also confused because, from what I can tell about your work above, are the x's in the numerator or the denominator?

OpenStudy (anonymous):

I believe the denominator (bottom)

OpenStudy (anonymous):

(1/7 * x^2 - 2/5)(1/2 *x + 1/2)

OpenStudy (anonymous):

thats how i read it....

OpenStudy (pfenn1):

\[\frac{1}{14x^3}+\frac{1}{14x^2}-\frac{1}{5x}-\frac{1}{5}\]

OpenStudy (anonymous):

then if i used the calculator how come it comes up with my answer that was shown ? how far off am I ? i am totallly confused now

OpenStudy (anonymous):

it says to find the product, is that not what I did?

OpenStudy (anonymous):

can you please show me the steps on how you got the answer ?

OpenStudy (pfenn1):

I just did the FOIL method.